-HAPTER 4 CLAUOVOR Tyrosine O Asparticauid Ho iX !! HOH Peptide Bond Alanie Y qoo Guysine Сн, / сн. н сH, Hн fan-e--c-coo-
Answer #1

The answers are explained in the images below.

is a terrapeptide The given peptide 4 residues amino acid a. It contains in the structure 2. There are 3 peptide bonds toontCH3 o ī 11 - c-NH-CH- C-NH-CH- (-NH-CH- ooo HN- CH-C-NM - H CH cooo -3 Charge = -1-1-1= coo CHH Cth + In HON- ctc-NECH -c-NEC